OpenStudy (anonymous):

Find the x-intercept of the line whose equation is 10x + 3y = 1. i got 1/3??

OpenStudy (saifoo.khan):

when x intercpts y = 0

OpenStudy (saifoo.khan):

10x + 3y = 1. 10(0) + 3y = 1. 3y = 1 y = 1/3

OpenStudy (saifoo.khan):

So u r correct.